OpenStudy (anonymous):

(y^2+6y-6)-(y^2+5y-8)

OpenStudy (anonymous):

you want to combine the like terms. so y^2 and y^2 should be combined the y's should be combined, and the plain number should be combined.

OpenStudy (anonymous):

but note that you're subtracting!!

OpenStudy (anonymous):

please show equation.

OpenStudy (anonymous):

(y^2-y^2) = 0 (6y-5y) = 1y = y (-6 - -8) = -6 + 8 = 2 so simplified, it would be 0 +y+2 = y+2